# Benjamin Carleton

### v6

@@ 2.3 Search Problems and Their Interreductions and Complexity
The definitions we just gave provide the tools we will use to show that the complexities of the two members of each known pair of collapsing standard types are polynomially related to each other (given oracle access to $W_{\mathcal{E}}$), and indeed that the members of the pair are closely related in terms of us being able to very efficiently build a solution to one from a solution to the other.666 Definitions 2.2 part 1 and 2.3 part 1 do not require that if $S$ is not a solution to $\mathcal{T}_{2}$ then the reduction declares that fact. Rather, the definitions are simply about efficiently obtaining a solution to $\mathcal{T}_{1}$ given a solution to $\mathcal{T}_{2}$. However, we mention that if one changed Definition 2.2 part 1 to require detection of nonsolution-hood, the set of pairs $(\mathcal{T}_{1},\mathcal{T}_{2})$ for which the reduction held would not change at all, since with the $W_{\mathcal{E}}$ oracle one can check whether $S$ is a solution to $\mathcal{T}_{2}$. Although Definition 2.3 part 1 is not in general guaranteed to be unchanged if it is altered to require detection of the case where $S$ is not a solution to $\mathcal{T}_{2}$, it clearly does remain unchanged by that alteration whenever $W_{\mathcal{E}}\in\mathrm{P}$. Most of the cases to which we apply Definition 2.3 indeed satisfy $W_{\mathcal{E}}\in\mathrm{P}$; in particular, plurality, veto, and approval voting each satisfy $W_{\mathcal{E}}\in\mathrm{P}$.
@@
We can now define what we mean by the search problem for $\mathcal{T}$ to be easy or hard. We say that the search problem for $\mathcal{T}$ is polynomial-time computable if there exists at least one refinement of $\widehat{\Pi}_{\mathcal{T}}$ (i.e., of $\widehat{\Pi}_{R_{\mathcal{T}}}$) that is a polynomial-time computable function. We say that the search problem for $\mathcal{T}$ is $\mathrm{NP}$-hard if for every refinement $f$ of $\widehat{\Pi}_{\mathcal{T}}$, it holds that $\mathrm{NP}\subseteq\mathrm{P}^{f}$. We say that the search problem for $\mathcal{T}$ is $\mathrm{NP}$-easy if there exists at least one refinement $f$ of $\widehat{\Pi}_{\mathcal{T}}$ that can be computed by a polynomial-time function given ${\rm SAT}$ as its oracle, i.e., is in the function class sometimes called “$\mathrm{FP}^{\mathrm{NP}}$” or “$\mathrm{PF}^{\mathrm{NP}}$.” If the search problem for $\mathcal{T}$ is both $\mathrm{NP}$-hard and $\mathrm{NP}$-easy, we say that the search problem is ${\rm SAT}$-equivalent. Given two SAT-equivalent search problems, this definition does not promise (as, in contrast, the second parts of Definition 2.3 and, in some sense, Definition 2.2 do) that one can produce a solution for one of the problems from any solution, for the same instance, to the other, and vice versa. However, the definitions just given do give us a clean way to make clear that one of our search problems is hard, or is easy.
@@ 3 Related Work
The papers most related to this one are those discussed already: Hemaspaandra, Hemaspaandra, and Menton [HHM20] and Carleton et al. [${}^{+}$22] collapsed existing control types as decision problems. And Carleton et al. [${}^{+}$22] posed as an open issue whether such collapses also collapse the associated search problems as to their complexity. Addressing that issue is the focus of the present paper.
@@
Control was created by Bartholdi, Tovey, and Trick [BTT92], and the 24 control types we study were developed in that paper and (for the introduction of destructive control types and stating the $\mathrm{TP}$/$\mathrm{TE}$ rules) in the work of Hemaspaandra, Hemaspaandra, and Rothe [HHR07]. The shift over time from the earliest papers’ focus on the $\mathrm{UW}$ model to a focus on either both $\mathrm{UW}$ and $\mathrm{NUW}$, or sometimes even just $\mathrm{NUW}$, is discussed, with citations, by Carleton et al. [${}^{+}$22, Related Work].
Many papers have studied control. Faliszewski and Rothe’s survey [FR16] is an excellent resource, with a rich range of citations to papers investigating the control complexities of specific election systems. Among the many systems that have been studied and seem to do somewhat well in resisting control attacks are Schulze and ranked pair elections [PX12, HLM16], Llull and Copeland elections [FHHR09, FHH15], normalized range voting [Men13], Bucklin and fallback elections [ER10, EPR11] (see also [EFRS15]), and SP-AV elections [ENR09].
Among the most popular election systems whose control complexity has been studied are approval [HHR07, ${}^{+}$10], plurality [BTT92, HHR07], and veto elections [Lin12, MR18, ERY19]. We will draw on some decision-problem algorithms from these papers to establish some of our search results. For example, Theorem 4.8 part 1 follows directly from the properties of an algorithm of Maushagen and Rothe [MR18].
@@ 4 Results
[…]
@@
$\mathcal{T}_{1}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{2}$: We give a polynomial-time algorithm performing the reduction, given access to oracle $W_{\mathcal{E}}$. On input $(I,S)$, where $I=(C,V,p)$ and $p\in C$, do the following. If the syntax is bad, halt. Otherwise, check that $S$ is a solution to $I$ under $\mathcal{T}_{2}$. This may involve up to three calls to the oracle. If $S$ is not a good solution, then simply halt as the definition does not require anything in this case (although note also the comments in Footnote 6). Since $S$ is a solution and our model is $\mathrm{NUW}$, $p$ must have participated and lost either in one of the two first-round elections or (in fact, exclusive-or) in the final-round election. In that election, let $C^{\prime}$ be the candidate set (note that $p\in C^{\prime}$ and $p$ is not a winner of $(C^{\prime},V)$). Then output as the successful $\mathcal{T}_{1}$ solution $C_{1}=C^{\prime}$ and $C_{2}=C-C^{\prime}$. In $\mathcal{T}_{1}$ on input $I$, $p$ will be eliminated in the $(C^{\prime},V)$ first-round election. [^7] This case shows why it is important that our definition is making an oracle to the winner problem of $\mathcal{E}$ available. Suppose we tried to claim that this direction held without any oracle use by, if $S$ is $(C_{1},C_{2})$, just outputting $S^{\prime}=(C_{1},C_{2})$. But then if $p$ participated in and lost in the $(C_{2},V)$ first-round $\mathcal{T}_{2}$ case, $S^{\prime}$ might not be a solution with respect to $\mathcal{T}_{1}$. Can we fix that by, if $p\in C_{2}$, just outputting $(C_{2},C_{1})$? No. Maybe in that case under $\mathcal{T}_{2}$ candidate $p$ lost in $(C_{2},V)$, or maybe it lost in $(\textrm{Winners}_{\mathcal{E}}(C_{2},V)\cup\textrm{Winners}_{\mathcal{E}}(C_{1},V),V)$. But if it was the latter, in $\mathcal{T}_{1}$ our second-round election is $(\textrm{Winners}_{\mathcal{E}}(C_{2},V)\cup C_{1},V)$ and it is possible that $p$ wins in that.
$\mathcal{T}_{2}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{1}$: We give the algorithm. On input $(I,S)$, where $I=(C,V,p)$ and $S=(C_{1},C_{2})$, check for bad syntax and by using the $W_{\mathcal{E}}$ oracle up to two times as per the nature of $\mathcal{T}_{1}$, check that $S$ is a solution to $I$ under $\mathcal{T}_{1}$. If the syntax is bad or $S$ is not a solution under $\mathcal{T}_{1}$, then simply halt. Otherwise, since $S$ is a solution under $\mathcal{T}_{1}$, $p$ was a participant in and eliminated either in the $(C_{1},V)$ contest or the $(C_{2}\cup\textrm{Winners}_{\mathcal{E}}(C_{1},V),V)$ contest. In the former case, output $(C_{1},C_{2})$, and in the latter case let $D=C_{2}\cup\textrm{Winners}_{\mathcal{E}}(C_{1},V)$ and output $(D,C-D)$. This is a solution for $\mathcal{T}_{2}$.
@@
Proof.    Let $\mathcal{E}$ be an election system and let $\mathcal{T}_{1}$, $\mathcal{T}_{2}$, and $\mathcal{T}_{3}$ be pairwise $\mathcal{E}$-matched control types such that $\mathcal{T}_{1}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{2}$ via $f$ and $\mathcal{T}_{2}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{3}$ via $g$ (with $f$ and $g$ both running in polynomial time given oracle $W_{\mathcal{E}}$). We will show that $\mathcal{T}_{1}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{3}$. On input $(I,S)$, where $I=(C,V,p)$, if $S$ is a solution for $\mathcal{T}_{3}$ on input $I$, then on input $((C,V,p),S)$ $g$ outputs a solution $S^{\prime}$ for $\mathcal{T}_{2}$ on input $I$. Additionally, if $S^{\prime}$ is a solution for $\mathcal{T}_{2}$ on input $I$, then on input $((C,V,p),S^{\prime})$ $f$ outputs a solution $S^{\prime\prime}$ for $\mathcal{T}_{1}$ on input $I$. Thus if $S$ is a solution for $\mathcal{T}_{3}$ on input $I$, then by applying $g$ and then $f$ in the manner just described we obtain, running in polynomial time with oracle $W_{\mathcal{E}}$, a solution $S^{\prime\prime}$ for $\mathcal{T}_{1}$ on input $I$. Thus $\mathcal{T}_{1}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{3}$.
@@
[…]
@@
We now explore the system-specific collapses from Carleton et al. [${}^{+}$22] in the same order that they appear in that paper. We observe that the way they establish their new decision-problem collapses for veto, approval, and election systems satisfying Unique-WARP is so constructive that the proofs implicitly show, for each new collapsing decision-problem pair $(\mathcal{T}_{1},\mathcal{T}_{2})$ that they find, that $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are polynomially search-equivalent, thus giving us the following corollary.
Corollary (to the proofs of Thms./Cor. 3.2/3.6/3.9/3.12/3.14/3.16–3.18 of [${}^{+}$22]) 4.5
@@
    For each election system $\mathcal{E}$ that satisfies Unique-WARP, -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ and -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ are polynomially search-equivalent.888 One might expect here and in part 3 of this corollary the weaker conclusion “polynomially search-equivalent with respect to $\mathcal{E}$.” But in both these parts we mean and prove “polynomially search-equivalent.”
@@
    For each election system $\mathcal{E}$ that satisfies Unique-WARP, -$\mathcal{E}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ and -$\mathcal{E}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ are polynomially search-equivalent.
@@
Proof.    For all parts except 2 and 3 we have $W_{\mathcal{E}}\in\mathrm{P}$ so we can check if the $S$ of the input is a valid solution, and so in those parts we below assume that input always is a solution for $I=(C,V,p)$ of the problem on the right-hand side of the reduction. For parts 2 and 3 we cannot and do not make that assumption.
@@
    $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}\leq_{\textnormal{search}}^{p}\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$: The proof of Theorem 3.2 of Carleton et al. [${}^{+}$22] shows how to construct, given $I=(C,V,p)$ and a solution $S=(C_{1},C_{2})$ for $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, a solution for $I$ to $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, and we note that this construction can easily be done in polynomial time.

### v7

@@ 2.3 Search Problems and Their Interreductions and Complexity
The definitions we just gave provide the tools we will use to show that the complexities of the two members of each known pair of collapsing standard types are polynomially related to each other (given oracle access to $W_{\mathcal{E}}$), and indeed that the members of the pair are closely related in terms of us being able to very efficiently build a solution to one from a solution to the other. [^7] Definitions 2.2 part 1 and 2.3 part 1 do not require that if $S$ is not a solution to $\mathcal{T}_{2}$ then the reduction declares that fact. Rather, the definitions are simply about efficiently obtaining a solution to $\mathcal{T}_{1}$ given a solution to $\mathcal{T}_{2}$. However, we mention that if one changed Definition 2.2 part 1 to require detection of nonsolution-hood, the set of pairs $(\mathcal{T}_{1},\mathcal{T}_{2})$ for which the reduction held would not change at all, since with the $W_{\mathcal{E}}$ oracle one can check whether $S$ is a solution to $\mathcal{T}_{2}$. Although Definition 2.3 part 1 is not in general guaranteed to be unchanged if it is altered to require detection of the case where $S$ is not a solution to $\mathcal{T}_{2}$, it clearly does remain unchanged by that alteration whenever $W_{\mathcal{E}}\in\mathrm{P}$. Most of the cases to which we apply Definition 2.3 indeed satisfy $W_{\mathcal{E}}\in\mathrm{P}$; in particular, plurality, veto, and approval voting each satisfy $W_{\mathcal{E}}\in\mathrm{P}$.
@@
We can now define what we mean by the search problem for $\mathcal{T}$ to be easy or hard. We say that the search problem for $\mathcal{T}$ is polynomial-time computable if there exists at least one refinement of $\widehat{\Pi}_{\mathcal{T}}$ (i.e., of $\widehat{\Pi}_{R_{\mathcal{T}}}$) that is a polynomial-time computable function. We say that the search problem for $\mathcal{T}$ is $\mathrm{NP}$-hard if for every refinement $f$ of $\widehat{\Pi}_{\mathcal{T}}$, it holds that $\mathrm{NP}\subseteq\mathrm{P}^{f}$. We say that the search problem for $\mathcal{T}$ is $\mathrm{NP}$-easy if there exists at least one refinement $f$ of $\widehat{\Pi}_{\mathcal{T}}$ that can be computed by a polynomial-time function given $\rm SAT$} as its oracle, i.e., is in the function class sometimes called $\mathrm{FP}^{\mathrm{NP}}$” or $\mathrm{PF}^{\mathrm{NP}}$.” If the search problem for $\mathcal{T}$ is both $\mathrm{NP}$-hard and $\mathrm{NP}$-easy, we say that the search problem is $\rm SAT$}-equivalent. Given two SAT-equivalent search problems, this definition does not promise (as, in contrast, the second parts of Definition 2.3 and, in some sense, Definition 2.2 do) that one can produce a solution for one of the problems from any solution, for the same instance, to the other, and vice versa. However, the definitions just given do give us a clean way to make clear that one of our search problems is hard, or is easy.
@@ 3 Related Work
The papers most related to this one are those discussed already: Hemaspaandra, Hemaspaandra, and Menton [HHM20] and Carleton et al. [CCH+24] collapsed existing control types as decision problems. And Carleton et al. [CCH+24] posed as an open issue whether such collapses also collapse the associated search problems as to their complexity. Addressing that issue is the focus of the present paper.
@@
Control was created by Bartholdi, Tovey, and Trick [BTT92], and the 24 control types we study were developed in that paper and (for the introduction of destructive control types and stating the $\mathrm{TP}$$\mathrm{TE}$ rules) in the work of Hemaspaandra, Hemaspaandra, and Rothe [HHR07a]. The shift over time from the earliest papers’ focus on the $\mathrm{UW}$ model to a focus on either both $\mathrm{UW}$ and $\mathrm{NUW}$, or sometimes even just $\mathrm{NUW}$, is discussed, with citations, by Carleton et al. [CCH+24, Related Work].
Many papers have studied control. Faliszewski and Rothe’s survey [FR16] is an excellent resource, with a rich range of citations to papers investigating the control complexities of specific election systems. Among the many systems that have been studied and seem to do somewhat well in resisting control attacks are Schulze and ranked pair elections [PX12, HLM16], Llull and Copeland elections [FHHR09, FHH15], normalized range voting [Men13], Bucklin and fallback elections [ER10, EPR11] (see also [EFRS15]), and SP-AV elections [ENR09]. Election control, in many variants, continues to be an area of active interest [GRSZ22, Yan22, Yan23, CGH+23, KRT23, KR24b, AHIJ+24, KR24a, MNN+24, Yan24, KR25].
Among the most popular election systems whose control complexity has been studied are approval [HHR07a, BEH+10], plurality [BTT92, HHR07a], and veto elections [Lin12, MR18, ERY19]. We will draw on some decision-problem algorithms from these papers to establish some of our search results. For example, Theorem 4.8 part 1 follows directly from the properties of an algorithm of Maushagen and Rothe [MR18].
@@ 4 Results
4.1 Search Equivalences
In the theorems and proofs that follow, $\textrm{Winners}_{\mathcal{E}}(C,V)$ will denote the set of winners of the election $(C,V)$ under election system $\mathcal{E}$. $\textrm{UniqueWinnerIfAny}_{\mathcal{E}}(C,V)$ will be $\textrm{Winners}_{\mathcal{E}}(C,V)$ if the latter has cardinality one, and will be the empty set otherwise. To compute these sets, we will typically leverage access to (in settings where one has such access) the oracle, $W_{\mathcal{E}}=\{(C,V,p)\mid p\in C$ and $p$ is a winner of the $\mathcal{E}$ election (C,V)}(C,V)\}, since clearly each of the new functions can be computed in polynomial time given access to $W_{\mathcal{E}}$.
@@ 4.1 Search Equivalences
$\mathcal{T}_{1}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{2}$: We give a polynomial-time algorithm performing the reduction, given access to oracle $W_{\mathcal{E}}$. On input $(I,S)$, where $I=(C,V,p)$ and $p\in C$, do the following. If the syntax is bad, halt. Otherwise, check that $S$ is a solution to $I$ under $\mathcal{T}_{2}$. This may involve up to three calls to the oracle. If $S$ is not a good solution, then simply halt as the definition does not require anything in this case (although note also the comments in Footnote 7). Since $S$ is a solution and our model is $\mathrm{NUW}$, $p$ must have participated and lost either in one of the two first-round elections or (in fact, exclusive-or) in the final-round election. In that election, let $C^{\prime}$ be the candidate set (note that $p\in C^{\prime}$ and $p$ is not a winner of $(C^{\prime},V)$). Then output as the successful $\mathcal{T}_{1}$ solution $C_{1}=C^{\prime}$ and $C_{2}=C-C^{\prime}$. In $\mathcal{T}_{1}$ on input $I$, $p$ will be eliminated in the $(C^{\prime},V)$ first-round election. [^8] This case shows why it is important that our definition is making an oracle to the winner problem of $\mathcal{E}$ available. Suppose we tried to claim that this direction held without any oracle use by, if $S$ is $(C_{1},C_{2})$, just outputting $S^{\prime}=(C_{1},C_{2})$. But then if $p$ participated in and lost in the $(C_{2},V)$ first-round $\mathcal{T}_{2}$ case, $S^{\prime}$ might not be a solution with respect to $\mathcal{T}_{1}$. Can we fix that by, if $p\in C_{2}$, just outputting $(C_{2},C_{1})$? No. Maybe in that case under $\mathcal{T}_{2}$ candidate $p$ lost in $(C_{2},V)$, or maybe it lost in $(\textrm{Winners}_{\mathcal{E}}(C_{2},V)\cup\textrm{Winners}_{\mathcal{E}}(C_{1},V),V)$. But if it was the latter, in $\mathcal{T}_{1}$ our second-round election is $(\textrm{Winners}_{\mathcal{E}}(C_{2},V)\cup C_{1},V)$ and it is possible that $p$ wins in that.
$\mathcal{T}_{2}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{1}$: We give the algorithm. On input $(I,S)$, where $I=(C,V,p)$ and $S=(C_{1},C_{2})$, check for bad syntax and by using the $W_{\mathcal{E}}$ oracle up to two times as per the nature of $\mathcal{T}_{1}$, check that $S$ is a solution to $I$ under $\mathcal{T}_{1}$. If the syntax is bad or $S$ is not a solution under $\mathcal{T}_{1}$, then simply halt. Otherwise, since $S$ is a solution under $\mathcal{T}_{1}$, $p$ was a participant in and eliminated either in the $(C_{1},V)$ contest or the $(C_{2}\cup\textrm{Winners}_{\mathcal{E}}(C_{1},V),V)$ contest. In the former case, output $(C_{1},C_{2})$, and in the latter case let $D=C_{2}\cup\textrm{Winners}_{\mathcal{E}}(C_{1},V)$ and output (D,CD)(D,C-D). This is a solution for $\mathcal{T}_{2}$. ❑
@@
Proof.  Let $\mathcal{E}$ be an election system and let $\mathcal{T}_{1}$, $\mathcal{T}_{2}$, and $\mathcal{T}_{3}$ be pairwise $\mathcal{E}$-matched control types such that $\mathcal{T}_{1}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{2}$ via $f$ and $\mathcal{T}_{2}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{3}$ via $g$ (with $f$ and $g$ both running in polynomial time given oracle $W_{\mathcal{E}}$). We will show that $\mathcal{T}_{1}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{3}$. On input $(I,S)$, where $I=(C,V,p)$, if $S$ is a solution for $\mathcal{T}_{3}$ on input $I$, then on input $((C,V,p),S)$ $g$ outputs a solution $S^{\prime}$ for $\mathcal{T}_{2}$ on input $I$. Additionally, if $S^{\prime}$ is a solution for $\mathcal{T}_{2}$ on input $I$, then on input $((C,V,p),S^{\prime})$ $f$ outputs a solution $S^{\prime\prime}$ for $\mathcal{T}_{1}$ on input $I$. Thus if $S$ is a solution for $\mathcal{T}_{3}$ on input $I$, then by applying $g$ and then $f$ in the manner just described we obtain, running in polynomial time with oracle $W_{\mathcal{E}}$, a solution $S^{\prime\prime}$ for $\mathcal{T}_{1}$ on input $I$. Thus $\mathcal{T}_{1}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{T}_{3}$. ❑
@@
-$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$: Say we are for $I=(C,V,p)$ given a purported solution $S=(C_{1},C_{2})$ to -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$. If $S$ is not a successful solution, immediately reject. Otherwise, either (a) $p\in C_{1}$ but $p$ is not a unique winner of $(C_{1},V)$, or (b) $p\in\textrm{UniqueWinnerIfAny}_{\mathcal{E}}(C_{1},V)\cup C_{2}$ yet $p$ is not a winner of $(\textrm{UniqueWinnerIfAny}_{\mathcal{E}}(C_{1},V)\cup C_{2},V)$. Using our oracle, determine which of (a) or (b) holds (exactly one must hold if $S$ was a solution). If (a) holds, output $(C_{1},C_{2})$. This is then a successful solution on $I$ to -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$. If (b) holds, then set $D=\textrm{UniqueWinnerIfAny}_{\mathcal{E}}(C_{1},V)\cup C_{2}$ and output (D,CD)(D,C-D) and this is a successful solution of $I$ to -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, since we know that $p$ is not a winner of $(D,V)$, so it certainly is not a unique winner of $(D,V)$, and so in our -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$ first round $p$ participates and is eliminated.
-$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$: Say we are given for $I=(C,V,p)$ a purported solution $S=(C_{1},C_{2})$ for $I$ to -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$. If $S$ is not a solution, immediately reject. Otherwise, we know that (a) $p\in C_{1}$ and $p$ does not uniquely win in $(C_{1},V)$, exclusive-or (b) $p\in\textrm{UniqueWinnerIfAny}_{\mathcal{E}}(C_{1},V)\cup C_{2}$ yet $p$ is not a unique winner of $(\textrm{UniqueWinnerIfAny}_{\mathcal{E}}(C_{1},V)\cup C_{2},V)$. If (a) holds, output $(C_{1},C_{2})$. This is a successful solution of $I$ for -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$ as $p$ is eliminated in the first round. If (b) holds, set $D=\textrm{UniqueWinnerIfAny}_{\mathcal{E}}(C_{1},V)\cup C_{2}$ and output (D,CD)(D,C-D). This is a successful solution of $I$ for -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$ as $p$ will be eliminated in the first round.
-$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}\leq_{\textnormal{search}}^{p,\mathcal{E}}\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$: Say we are given $I=(C,V,p)$ and a purported solution $S=(C_{1},C_{2})$ for -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ for $I$. If $S$ is not a solution, immediately reject. Otherwise, since $C_{1}$ and $C_{2}$ are symmetric in $\mathrm{RPC}$, w.l.o.g. assume $p\in C_{1}$ (otherwise, if $p\in C_{2}$, rename $C_{1}$ and $C_{2}$ so that $p\in C_{1}$). So exactly one of (a) and (b) holds, where (a) and (b) are: (a) $p\in C_{1}$ and $p$ is not a unique winner of $(C_{1},V)$, and (b) $p\in\textrm{UniqueWinnerIfAny}_{\mathcal{E}}(C_{1},V)\cup\textrm{UniqueWinnerIfAny}_{\mathcal{E}}(C_{2},V)$ and $p$ is not a unique winner of $(\textrm{UniqueWinnerIfAny}_{\mathcal{E}}(C_{1},V)\cup\textrm{UniqueWinnerIfAny}_{\mathcal{E}}(C_{2},V),V)$. If (a) holds, output $(C_{1},C_{2})$ and that is a successful solution for $I$ of -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ as $p$ is eliminated in the first round. If (b) holds, set $D=\textrm{UniqueWinnerIfAny}_{\mathcal{E}}(C_{1},V)\cup\textrm{UniqueWinnerIfAny}_{\mathcal{E}}(C_{2},V)$, and output (D,CD)(D,C-D) and that is a successful solution for $I$ of -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ as $p$ is eliminated in the first round. ❑
@@
We now explore the system-specific collapses from Carleton et al. [CCH+24] in the same order that they appear in that paper. We observe that the way they establish their new decision-problem collapses for veto, approval, and election systems satisfying Unique-$\alpha$ is so constructive that the proofs implicitly show, for each new collapsing decision-problem pair $(\mathcal{T}_{1},\mathcal{T}_{2})$ that they find, that $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ are polynomially search-equivalent, thus giving us the following corollary.
Corollary (to the proofs of Thms./Cor. 3.2/3.6/3.9/3.12/3.14/3.16–3.18 of [CCH+24]) 4.5
@@
    For each election system $\mathcal{E}$ that satisfies Unique-$\alpha$, -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ and -$\mathcal{E}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$ are polynomially search-equivalent. [^9] One might expect here and in part 3 of this corollary the weaker conclusion “polynomially search-equivalent with respect to $\mathcal{E}$.” But in both these parts we mean and prove “polynomially search-equivalent.”
@@
    For each strongly voiced election system $\mathcal{E}$ that satisfies Property $\alpha$, -$\mathcal{E}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\mathrm{PC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ and -$\mathcal{E}\hbox{-}\allowbreak\mathrm{CC}\hbox{-}\mathrm{RPC}\hbox{-}\allowbreak\mathrm{TP}\hbox{-}\allowbreak\mathrm{UW}$ are polynomially search-equivalent.
@@
Proof.  For all parts except 2 and 3, we have $W_{\mathcal{E}}\in\mathrm{P}$ so we can check if the $S$ of the input is a valid solution, and so in those parts we below assume that input always is a solution for $I=(C,V,p)$ of the problem on the right-hand side of the reduction. For parts 2 and 3, we cannot and do not make that assumption.
@@
    $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}\leq_{\textnormal{search}}^{p}\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$: The proof of Theorem 3.2 of Carleton et al. [CCH+24] shows how to construct, given $I=(C,V,p)$ and a solution $S=(C_{1},C_{2})$ for $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{UW}$, a solution for $I$ to $\mathrm{Veto}\hbox{-}\allowbreak\mathrm{DC}\hbox{-}\allowbreak\mathrm{PV}\hbox{-}\allowbreak\mathrm{TE}\hbox{-}\allowbreak\mathrm{NUW}$, and we note that this construction can easily be done in polynomial time.